Assume that (G,*) is a cyclic group and that H is a subgroup of G. Explain why H is a normal subgroup G,and prove that G/H is cyclic.
1
Expert's answer
2020-11-24T17:31:21-0500

Each cyclic group "G=\langle a \rangle" is Abelian. Indeed, "a^ma^n=a^{m+n}=a^{n+m}=a^na^m". Therefore, "gH=Hg" for all "g\in G", and "H" is a normal subgroup of "G".


Let "[g]=gH" be arbitrary element of "G\/H". Since "G=\langle a \rangle" is cyclic, "g=a^k" for some "k\in \mathbb Z." By defenition of product of elements of "G\/H" , "[g]=[a^k]=[a]^k." Therefore, "G/H=\langle[a]\rangle," and "G\/H"

is cyclic.